# Second-Order Constrained Dynamic Optimization

### v2

@@ A.1.2 Merit function and line search in SQP
To determine an appropriate step size $\alpha$, we use Augmented Lagrangian (AL) merit function and achieve constraint satisfaction and cost reduction as in Gill et al. (1986b). With the AL merit function, the appropriate step size is determined by line search after setting the nonnegative penalty parameter $\rho\in\mathbb{R}^{w}$ so that it ensures the existence of good $\alpha$. Let $\phi$ be the AL merit function. $\phi$ is defined as a function of the penalty parameter parameterized by $\alpha$ as
@@
$\quad-(\lambda_{k}+\alpha\delta\lambda_{k})^{\mathsf{T}}[c(x_{k}+\alpha\delta x_{k})-(s+\alpha\delta s)]$
$\quad\quad+\sum_{i=1}^{w}\frac{1}{2}\rho_{i}\big{[}c_{i}(x_{k}+\alpha\delta x_{k})-[s+\alpha\delta s]_{i}\big{]}^{2},$
where $s\in\mathbb{R}^{w}$ is a slack variable that is introduced only for AL and line search and thus has nothing to do with the QP subproblem. $\delta\lambda_{k}$ here is defined as a difference between optimal $\lambda_{k}$ denoted by $\lambda^{\ast}_{k}$ and current $\lambda_{k}$, that is
@@
$\lambda^{\ast}_{k}$ is obtained as byproducts of the solution of (66). $\delta x_{k}$ is an optimal solution for (66), but we drop $\ast$ to maintain readability. The slack variable $s$ is initialized by
$s_{i}=\begin{cases}\max\{0,c_{i}(x_{k})\},&\rho_{i}=0,\\ \max\{0,c_{i}(x_{k})-\lambda_{i}/\rho_{i}\},&\text{otherwise}.\\ \end{cases}$
@@
$c(x_{k})+\nabla c(x_{k})\delta x_{k}=s+\delta s.$ (70)
Since only the current iteration’s variables matter in the merit function, we drop subscript $k$ for iteration hereafter. To ensure the existence of $\alpha$, $\rho$ needs to make $\phi^{\prime}(\rho_{k};\alpha=0)$ sufficiently negative, where $\phi^{\prime}=\frac{\partial\phi}{\partial\alpha}$. This condition is typically given as
@@
$\text{with}\ \quad a_{i}$ $=[c_{i}(x)-s_{i}]^{2},$
$b=[\nabla f_{0}(x)]^{\mathsf{T}}\delta x$ $+\frac{1}{2}\delta x^{\mathsf{T}}H\delta x-[\delta\lambda-\lambda]^{\mathsf{T}}[c(x)-s].$
@@
The left-hand side of the constraint is computed as
@@
$-\delta\lambda^{\mathsf{T}}\big{[}c(x)-s\big{]}-\lambda^{\mathsf{T}}[\nabla c(x)\delta x-s]$
$+\sum_{i}^{w}\rho_{i}[c_{i}(x)-s_{i}][\nabla c_{i}(x)\delta x-\delta s_{i}].$
@@
$[\nabla f_{0}(x)]^{\mathsf{T}}\delta x+\frac{1}{2}\delta x^{\mathsf{T}}H\delta x-[\delta\lambda-\lambda]^{\mathsf{T}}[c(x)-s]$ (75)
$\leq\sum_{i=1}^{w}\rho_{i}[c_{i}(x)-s_{i}]^{2}.$
Notice that the first two terms are the optimal value of QP subproblem (68). Now, the problem given in (73) is written as
[…]
When the left-hand side of the inequality constraint (75) is not positive, the solution for (76) is clearly $\rho_{i}=0,\forall i=1,\cdots w$. We now consider the case where the left-hand side of the constraint is positive. To simplify the expression, we use $a$, $b$ in (71) and write (76) as
@@
where $b>0$ is because we are considering the case where the left-hand side of the constraints in (76) to be positive. Lagrangian of this problem is given by
@@
with corresponding multipliers $\lambda_{1}\in\mathbb{R}$ and $\lambda_{2}\in\mathbb{R}^{w}$. KKT condition gives
$\rho-{\lambda_{1}}a-\lambda_{2}$ $=0$ (78a)
${\lambda_{1}}\geq 0,\quad\lambda_{1}(a^{\mathsf{T}}\rho-b)$ $=0$ (78b)
$\lambda_{2,i}\geq 0,\quad\lambda_{2,i}\rho_{i}$ $=0,\quad i=1\cdots w$ (78c)
@@
If $\lambda_{1}=0$, $\rho=0$ from (79), causing $a^{\mathsf{T}}\rho=0$, again. Therefore, $\lambda_{1}\neq 0$. This gives
$\lambda_{1}=\frac{b}{||a||^{2}}.$
@@ A.1.3 Hessian Update
After updating $x$ with appropriate $\alpha$, the gradient and Hessian of the QP subproblem are updated for the next iteration. The gradient is updated by computing the gradient with the new $x$. For the Hessian, since its exact computation is expensive and since it needs to be PD, BFGS quasi-Newton update (Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970) is used, which is give as follows.
@@
$y_{k}$ $=\nabla\mathcal{L}_{\rm{m}}(x_{k+1},x_{k},\lambda_{k+1})-\nabla\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k+1})$ (85)
@@
$\hskip 22.76219pt-[\nabla c(x_{k+1})-\nabla c(x_{k})]^{\mathsf{T}}\lambda_{k+1},$
@@
for an objective $f_{0}(x)\in\mathbb{R}$, constant $c$ and a search direction $p_{k}$ to make a successful progress (Nocedal and Wright, 2006). In our setting, the left-hand side is $\phi^{\prime}(\alpha)$ and the right-hand side is a constant times $\phi^{\prime}(0)$. This condition was originally used to decide when to terminate the line search. Setting $c=1$ gives the condition used here. When $y_{k}^{\mathsf{T}}\delta x_{k}$ is not positive or has a very small positive value, the update law is modified to ensure that $H_{k+1}$ is PD. The modification is invoked when
@@
$\mathcal{L}_{\rm{mA}}(x,x_{k},\lambda_{k};\omega)$ $=f_{0}(x)-\lambda_{k}^{\mathsf{T}}d_{\rm{L}}(x,x_{k})$
@@
The third penalty term with the penalty parameter $\omega$ is added to the modified Lagrangian in (67). Although the names are similar, this $\mathcal{L}_{\rm{mA}}$ has nothing to do with the AL merit function $\phi$ in (69) used to find the step size $\alpha$. $\mathcal{L}_{\rm{mA}}$ is defined only to modify $y_{k}$ to update Hessian. Here, new $y_{k}$ is obtained as a sum of current $y_{k}$ and difference $\delta y_{k}$ as
@@
$\hskip 34.1433pt+\underbrace{[\nabla c(x_{k+1})-\nabla c(x_{k})]^{\mathsf{T}}\Omega d_{\rm{L}}(x_{k+1},x_{k})}_{\delta y_{k}}.$
@@
$=\big{[}[\nabla c(x_{k+1})-\nabla c(x_{k})]^{\mathsf{T}}\Omega d_{\rm{L}}(x_{k+1},x_{k})\big{]}^{\mathsf{T}}\delta_{k}$
$=\big{[}d_{\rm{L}}(x_{k+1},x_{k})\odot[\nabla c(x_{k+1})-\nabla c(x_{k})]\delta_{k}\big{]}^{\mathsf{T}}\omega.$
In order to set up a problem to find $\omega$, define $\bar{a}\in\mathbb{R}^{w}$ and $\bar{b}\in\mathbb{R}$ as
$\bar{a}$ $=d_{\rm{L}}(x_{k+1},x_{k})\odot[\nabla c(x_{k+1})-\nabla c(x_{k})]\delta_{k},$
@@
$\lambda_{2}||\bar{a}||^{2}=\bar{b}\Leftrightarrow\lambda_{2}=\frac{\bar{b}}{||\bar{a}||^{2}},$
@@ A.1.4 Merit function in SQP for dynamical systems
This section show show the AL merit function can determine the step size $\alpha$ in single- and multiple-shooting SQP for dynamical systems. We take the multiple-shooting formulation as an example. Consider an optimization problem of cost $J$ associated with the trajectory of the dynamical system under constraints.
[…]
$\hskip 17.07164ptc_{N}(x_{N})\geq 0,$
where $c_{k}(x_{k},u_{k})\in\mathbb{R}^{w}$, and a corresponding multiple shooting SQP
[…]
@@
$\text{with}\ C_{\alpha}(\delta Y)=C(\bar{Y}+\alpha\delta Y),\ \hat{F}_{\alpha}(\delta Y)=\hat{F}(\bar{Y}+\alpha\delta Y).$
@@
The minimum norm $\rho_{\rm{v}}$ that satisfies $\phi(\rho_{\rm{v}},0)^{\prime}\leq-\frac{1}{2}\delta Y^{\mathsf{T}}[\nabla_{YY}\mathcal{L}]\delta Y$ is obtained by substituting $\hat{a}_{2}$ and $\hat{b}$ for $a$ and $b$ in (72), followed by the dumping mechanism in (81). With this penalty parameter, the line search is performed as in section A.1.2. In practice, we set a minimum value of $\alpha$ in the line search. When a reasonable step size is not found, we discard the solution and regularize the Hessian as in (35), solving the QP subproblem again with the regularized Hessian.
@@ A.1.5 Hessian Update
$\mathcal{L}_{\rm{m}}(Y,Y_{k},\lambda_{\rm{v}},\nu_{\rm{v}})$ $=J(Y)-\lambda_{\rm{v}}^{\mathsf{T}}d_{\rm{LI}}(Y,Y_{k})-\nu_{\rm{v}}^{\mathsf{T}}d_{\rm{LE}}(Y,Y_{k}),$
@@
If the curvature is not large enough, we utilize modified AL
@@
$=$ $J(Y)-\lambda_{\rm{v}}^{\mathsf{T}}d_{\rm{LI}}(Y,Y_{k})+\frac{1}{2}d_{\rm{LI}}(Y,Y_{k})^{\mathsf{T}}\Omega_{\rm{I}}d_{\rm{LI}}(Y,Y_{k})$
$\hskip 22.76219pt-\nu_{\rm{v}}^{\mathsf{T}}d_{\rm{LE}}(Y,Y_{k})+\frac{1}{2}d_{\rm{LE}}(Y,Y_{k})^{\mathsf{T}}\Omega_{\rm{E}}d_{\rm{LE}}(Y,Y_{k}),$
@@ B.2 Multipliers in the recursion
Using these $\hat{Q}^{\dagger}$, they update $V$ with gains of $u_{k}$, i.e., $\kappa$ and $K$ in (36). Observe that the gains for $\lambda_{k}$, that is, $r$ and $R$ are not used. This can be seen as taking $\delta\lambda_{k}\approx 0$ during the recursion of the value function, which may not be true. The authors also remove the constraints at the terminal time step. In most cases, the optimal solution is located within the feasible region. Hence, the modification would not significantly affect the problem. However, we have observed that the feedback gain of $\lambda_{N}$, increases when the corresponding $s_{N}$ approaches zero. As a result, the fraction to the boundary rule for DDP in (51) is not satisfied even with a small step size $\alpha$ because it can only affect the feedforward part during the line search. The numerical instability of $s^{-1}$ when $s$ becomes close to zero is a typical problem with the IP method. One measure is the change of variables by multiplying $S$ to remove the term $s^{-1}$ (Nocedal and Wright, 2006), but this is not implemented in IP DDP.
In the experiment in section 7, we used the same formulation as in the original work, i.e., $\delta\lambda\approx 0$ in the recursion of the value function and remove the constraints at the final time step.
@@ D.3.1 Dynamics and Parameters
$M(q){\ddot{q}}$
$=$ $F+[m_{p}L\dot{\phi}^{2}\sin{\phi},-(m_{p}+m_{q})g_{0}-m_{p}L\dot{\phi}^{2}\cos{\phi},$
$\hskip 42.67912pt0,-m_{p}Lg_{0}\sin{\phi}]^{\mathsf{T}},$
@@ D.4 Swimmer
D.5 Additional Results for ADMM DDP

### v3

@@ A.1.2 Merit function and line search in SQP
To determine an appropriate step size $\alpha$, we use Augmented Lagrangian (AL) merit function that achieves constraint satisfaction and cost reduction (Gill et al., 1986b). With the AL merit function, the appropriate step size is found by line search after setting the nonnegative penalty parameter $\rho\in\mathbb{R}^{w}$ so that it ensures the existence of good $\alpha$. Let $\phi$ be the AL merit function. $\phi$ is defined as a function of the penalty parameter parameterized by $\alpha$ as
@@
$\quad-(\lambda_{k}+\alpha\delta\lambda_{k})^{\mathsf{T}}[g(x_{k}+\alpha\delta x_{k})-(s+\alpha\delta s)]$
$\quad\quad+\sum_{i=1}^{w}\frac{1}{2}\rho_{i}\big{[}g_{i}(x_{k}+\alpha\delta x_{k})-[s+\alpha\delta s]_{i}\big{]}^{2},$
where $s\in\mathbb{R}^{w}$ is a slack variable that is introduced only for AL and line search. $\delta\lambda_{k}$ here is defined as a difference between optimal $\lambda_{k}$ denoted by $\lambda^{\ast}_{k}$ and current $\lambda_{k}$, that is
@@
$\lambda^{\ast}_{k}$ is obtained as byproducts of the solution of (25). Here, we respect the original work and have negative sign for $\lambda$. $\delta x_{k}$ is the solution for (25), but we drop $\ast$ for readability. The slack variable $s$ is initialized by
$s_{i}=\begin{cases}\max\{0,g_{i}(x_{k})\},&\rho_{i}=0,\\ \max\{0,g_{i}(x_{k})-\lambda_{i}/\rho_{i}\},&\text{otherwise}.\\ \end{cases}$
@@
$g(x_{k})+\nabla g(x_{k})\delta x_{k}=s+\delta s.$ (69)
Since only the current iteration’s variables matter in the merit function, we drop subscript $k$ for iteration hereafter. To ensure the existence of $\alpha$, $\rho$ needs to make $\phi^{\prime}(\rho_{k};\alpha=0)$ sufficiently negative, where $\phi^{\prime}=\frac{\partial\phi}{\partial\alpha}$. This condition is typically given by
@@
$\text{with}\ \quad a_{i}$ $=[g_{i}(x)-s_{i}]^{2},$
$b=[\nabla f_{0}(x)]^{\mathsf{T}}\delta x$ $+\frac{1}{2}\delta x^{\mathsf{T}}H\delta x-[\delta\lambda-\lambda]^{\mathsf{T}}[g(x)-s].$
@@
The LHS of the constraint is computed as
@@
$-\delta\lambda^{\mathsf{T}}\big{[}g(x)-s\big{]}-\lambda^{\mathsf{T}}[\nabla g(x)\delta x-s]$
$+\sum_{i}^{w}\rho_{i}[g_{i}(x)-s_{i}][\nabla g_{i}(x)\delta x-\delta s_{i}].$
@@
$[\nabla f_{0}(x)]^{\mathsf{T}}\delta x+\frac{1}{2}\delta x^{\mathsf{T}}H\delta x-[\delta\lambda-\lambda]^{\mathsf{T}}[g(x)-s]$ (74)
$\leq\sum_{i=1}^{w}\rho_{i}[g_{i}(x)-s_{i}]^{2}.$
Notice that the first two terms are the optimal value of QP subproblem (67). Now, the constraint of the problem (72) is equivalent to (74). When the LHS of (74) is not positive, the solution for (72) is $\rho_{i}=0,\forall i=1,\cdots w$. We consider the case where the LHS of the constraint is positive. To simplify the expression, we use $a$, $b$ in (71) and write (72) as
@@
where $b>0$ is because we are considering the case where the LHS of the constraints in (74) is positive. Lagrangian of this problem is given by
@@
with corresponding multipliers $\lambda_{1}\in\mathbb{R}$ and $\lambda_{2}\in\mathbb{R}^{w}$. KKT condition yields
$\rho-{\lambda_{1}}a-\lambda_{2}$ $=0,$ (76a)
${\lambda_{1}}\geq 0,\quad\lambda_{1}(a^{\mathsf{T}}\rho-b)$ $=0,$ (76b)
$\lambda_{2,i}\geq 0,\quad\lambda_{2,i}\rho_{i}$ $=0,\quad i=1\cdots w,$ (76c)
@@
If $\lambda_{1}=0$, $\rho=0$, from (77), we have $a^{\mathsf{T}}\rho=0$, again. Therefore, $\lambda_{1}\neq 0$, which gives
$\lambda_{1}=b/{||a||^{2}}.$
@@ A.1.3 Hessian Update
After updating $x$ with appropriate $\alpha$, the new subproblem is obtained, whose gradient is computed with the new $x$. For the Hessian, since its exact computation is expensive and needs to be PD, BFGS quasi-Newton update (Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970) is used, which is give as follows.
@@
$y_{k}$ $=\nabla\mathcal{L}_{\rm{m}}(x_{k+1},x_{k},\lambda_{k+1})-\nabla\mathcal{L}_{\rm{m}}(x_{k},x_{k},\lambda_{k+1})$
@@
$\hskip 22.76219pt+[\nabla g(x_{k+1})-\nabla g(x_{k})]^{\mathsf{T}}\lambda_{k+1},$
@@
for an objective $f_{0}(x)\in\mathbb{R}$, constant $c$ and a search direction $p_{k}$ to make a successful progress (Nocedal and Wright, 2006). In our setting, the LH side is $\phi^{\prime}(\alpha)$ and the RH side is a constant times $\phi^{\prime}(0)$. This condition was originally used to decide when to terminate the line search. Setting $c=1$ gives the condition used here. When $y_{k}^{\mathsf{T}}\delta x_{k}$ is not positive or has a very small positive value, the update law is modified to ensure that $H_{k+1}$ is PD. The modification is invoked when
@@
$\mathcal{L}_{\rm{mA}}(x,x_{k},\lambda_{k};\omega)$ $=f_{0}(x)+\lambda_{k}^{\mathsf{T}}d_{\rm{L}}(x,x_{k})$
@@
The third penalty term with the parameter $\omega$ is added to the modified Lagrangian in (66). Although the names are similar, this $\mathcal{L}_{\rm{mA}}$ has nothing to do with the AL merit function $\phi$ in (68) used to find the step size $\alpha$. $\mathcal{L}_{\rm{mA}}$ is defined only to modify $y_{k}$ to update Hessian. Here, new $y_{k}$ is obtained as a sum of current $y_{k}$ and difference $\delta y_{k}$ as
@@
$\hskip 34.1433pt+\underbrace{[\nabla g(x_{k+1})-\nabla g(x_{k})]^{\mathsf{T}}\Omega d_{\rm{L}}(x_{k+1},x_{k})}_{\delta y_{k}}.$
@@
$=\big{[}[\nabla g(x_{k+1})-\nabla g(x_{k})]^{\mathsf{T}}\Omega d_{\rm{L}}(x_{k+1},x_{k})\big{]}^{\mathsf{T}}\delta_{k}$
$=\big{[}d_{\rm{L}}(x_{k+1},x_{k})\odot[\nabla g(x_{k+1})-\nabla g(x_{k})]\delta_{k}\big{]}^{\mathsf{T}}\omega.$
To set up a problem to find $\omega$, define $\bar{a}\in\mathbb{R}^{w}$ and $\bar{b}\in\mathbb{R}$ as
$\bar{a}$ $=d_{\rm{L}}(x_{k+1},x_{k})\odot[\nabla g(x_{k+1})-\nabla g(x_{k})]\delta_{k},$
@@
$\lambda_{2}||\bar{a}||^{2}=\bar{b}\Leftrightarrow\lambda_{2}=\bar{b}/||\bar{a}||^{2},$
@@ A.1.4 Merit function in SQP for dynamical systems
This section show how the AL merit function can determine the step size $\alpha$ in SQP for dynamical systems. We take the multiple-shooting formulation as an example. Consider an optimization problem in (64).
@@
$\text{with}\ G_{\alpha}(\delta Y)=G(\bar{Y}+\alpha\delta Y),\ \hat{F}_{\alpha}(\delta Y)=\hat{F}(\bar{Y}+\alpha\delta Y).$
@@
The minimum norm $\rho_{\rm{v}}$ that satisfies $\phi(\rho_{\rm{v}},0)^{\prime}\leq-\frac{1}{2}\delta Y^{\mathsf{T}}[\nabla_{YY}\mathcal{L}]\delta Y$ is obtained by substituting $\hat{a}_{2}$ and $\hat{b}$ for $a$ and $b$ in (71), followed by the dumping mechanism in (79). With this penalty parameter, the line search is performed as in section A.1.2. When a reasonable step size is not found with a small $\alpha$, we discard the solution and regularize the Hessian as in (35), solving the QP subproblem again with the regularized Hessian.
@@ A.1.5 Hessian Update
$\mathcal{L}_{\rm{m}}(Y,Y_{k},\lambda_{\rm{v}},\nu_{\rm{v}})$ $=J(Y)+\lambda_{\rm{v}}^{\mathsf{T}}d_{\rm{LI}}(Y,Y_{k})+\nu_{\rm{v}}^{\mathsf{T}}d_{\rm{LE}}(Y,Y_{k}),$
@@
If the curvature is not large enough, we use modified AL
@@
$=$ $J(Y)+\lambda_{\rm{v}}^{\mathsf{T}}d_{\rm{LI}}(Y,Y_{k})+\frac{1}{2}d_{\rm{LI}}(Y,Y_{k})^{\mathsf{T}}\Omega_{\rm{I}}d_{\rm{LI}}(Y,Y_{k})$
$\hskip 22.76219pt+\nu_{\rm{v}}^{\mathsf{T}}d_{\rm{LE}}(Y,Y_{k})+\frac{1}{2}d_{\rm{LE}}(Y,Y_{k})^{\mathsf{T}}\Omega_{\rm{E}}d_{\rm{LE}}(Y,Y_{k}),$
@@ B.2 Multipliers in the recursion
Using these $\hat{Q}^{\dagger}$, they update $V$ with gains of $u_{k}$, i.e., $\kappa$ and $K$ in (36). Observe that the gains for $\lambda_{k}$, that is, $r$ and $R$ are not used. This can be seen as taking $\delta\lambda_{k}\approx 0$ during the recursion of the value function, which may not be true. The authors also remove the constraints at the terminal time step. In most cases, the optimal solution is located within the feasible region. Hence, the modification would not significantly affect the problem. However, we have observed that the feedback gain of $\lambda_{N}$ increases when the corresponding $s_{N}$ approaches zero. As a result, the fraction to the boundary rule for DDP in (51) is not satisfied even with a small step size $\alpha$ because it can only affect the feedforward part during the line search. The numerical instability of $s^{-1}$ when $s$ becomes close to zero is a typical problem with the IP method. One measure is the change of variables by multiplying $S$ to remove the term $s^{-1}$ (Nocedal and Wright, 2006), but this is not implemented in IP DDP. In the experiment in section 7, we used the same formulation as in the original work, i.e., $\delta\lambda\approx 0$ in the recursion of the value function and remove the constraints at the final time step.
@@ D.3.1 Dynamics and Parameters
$M(q){\ddot{q}}=F+[m_{p}L\dot{\phi}^{2}\sin{\phi},$
$\quad\ -(m_{p}+m_{q})g_{0}-m_{p}L\dot{\phi}^{2}\cos{\phi},0,-m_{p}Lg_{0}\sin{\phi}]^{\mathsf{T}},$
@@ D.4 Swimmer
D.5 Panda
This section provides the details of the experiment with Panda. We denote the position and velocity of the end effector as
$x_{\mathrm{e}}=[x_{\mathrm{ep}}^{\mathsf{T}},v_{\mathrm{e}}^{\mathsf{T}},\omega_{\mathrm{e}}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{R}^{12},$
where $x_{\mathrm{ep}}$, $v_{\mathrm{e}}$, and $\omega_{\mathrm{e}}$ are position, velocity and angular velocity, respectively. We set the targets for the tasks by specifying the position $x_{\mathrm{ep}}$ and setting $v_{\mathrm{e}}=\omega_{\mathrm{e}}=0$. The quantity $x_{\mathrm{e}}$ is computed from the state $x$. The limits of the joints are based on the suggested limits, and the command limits are from the xml file of the model Zakka et al. (2022). To encode obstacle constraints, we place spheres along the links of the arm so that the spheres cover the link. Then, we define constraints between the spheres and obstacles. The time horizon and discretization interval of the problem are $N=100$ and $\Delta t$ = 0.02. Cost matrices are as follows:
$R_{1}$ $=0.01I_{7},\ R_{2}=\mathrm{blkdiag}[0.8I_{3},0.2I_{3},0.02I_{3}]$
$Q$ $=\mathrm{blkdiag}[100I_{3},I_{3},0.1I_{3}],$
where $\mathrm{blkdiag}$ stands for block diagonal matrix.
D.6 Additional Results for ADMM DDP